The sum of the digits of a three-digit number is 12. The tens digit is two times the hundreds digit. The sum of the units and tens digits is three times the hundreds digit. Find the number.
Answer by mananth(16949)

Let the 3 digit numbber be xyz
x+y+z=12
y=2x
y+z= 3x

y=2x
substitute y =2x in
Y+z =3x
2x+z=3x
z=3x-2x
z=x
x+y+z=12
x+2x+x=12
4x=12
/4
x=3
z=x so z=3
x+y+z=12
3+y+3=12
y+6=12
y=12-6
y=6
so the number is 363